# OmniQuery: Contextually Augmenting Captured Multimodal Memory to Enable Personal Question Answering

### v1

@@ Abstract.
People often capture memories through photos, screenshots, and videos. While existing AI-based tools enable querying this data using natural language, they mostly only support retrieving individual pieces of information like certain objects in photos, and struggle with answering more complex queries that involve interpreting interconnected memories like event sequences. We conducted a one-month diary study to collect realistic user queries and generated a taxonomy of necessary contextual information for integrating with captured memories. We then introduce OmniQuery, a novel system that is able to answer complex personal memory-related questions that require extracting and inferring contextual information. OmniQuery augments single captured memories through integrating scattered contextual information from multiple interconnected memories, retrieves relevant memories, and uses a large language model (LLM) to comprehensive answers. In human evaluations, we show the effectiveness of OmniQuery with an accuracy of 71.5% and it outperformed a conventional RAG system, winning or tying in 74.5% of the time.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Natural language interfacesccs: Human-centered computing User studies
Figure 1. OmniQuery enables answering natural language personal questions (a) on individuals’ captured memories such as captured photos, saved screenshots and recorded videos (b). This is achieved by augmenting the captured memories through identifying and integrating scattered contextual information from multiple interconnected memories (c). Using the augmented contextual information, OmniQuery retrieves relevant memories, and leverages an LLM to generate a comprehensive answer and the reference memories (d).
@@ 1. INTRODUCTION
People often record their everyday life by taking photos, screenshots, and videos, whether for saving important information, documenting special occasions, or simply capturing a funny moment (Li et al., 2024b). These captured instances, referred to as captured memories, collectively represent subsets of an individual’s episodic memories (Tulving, 2002), a type of long-term memory that contains both specific past experiences and associated contextual details. These episodic memories are essential for answering memory-related personal questions like, ”What social events did I attend during CHI 2024?” (Figure 1a), which can help users reflect on past experiences and make informed decisions in daily tasks.
@@
However, these raw captured memories by themselves are insufficient to answer personal questions, as they lack contextual details that are typically implicit and scattered across multiple pieces of data. For examples, as shown in Figure 1b, memories of attending parties during CHI 2024 are not explicitly annotated as occurring during the event. Answering such personal questions requires extracting and integrating contextual information not typically contained within a single captured instance. For example, by integrating multiple memories that mention “CHI 2024” in their content and extracting their metadata, it is possible to determine when the users attended the conference and connect related social events memories from that period to CHI 2024 (Figure 1c), enabling the answer of the query (Figure 1d).
@@
Advancements in AI have enable question answering (QA) on long documents (Beltagy et al., 2020; Wang et al., 2024), knowledge graph (Huang et al., 2019a; Yasunaga et al., 2021a), multimodal databases (Talmor et al., 2021; Chen et al., 2022), egocentric videos (Mangalam et al., 2023; Hou et al., 2023). These methods typically rely on data-driven approaches to train powerful models for the target task. However, the private nature of captured memories makes it difficult to curate large datasets, posing challenges for training models specifically for QA on personal captured memories. Recent LLM-based work has adopted retrieval augmented generation (RAG) workflow to handle external databases without specific training (Lewis et al., 2021). However, such methods depend on explicit connections between queries and relevant external data (Edge et al., 2024). In contrast, captured memories are often unstructured and lack contextual annotations, making it difficult to establish explicit links between queries and interconnected memories.
@@
To achieve this, we propose OmniQuery, a novel approach designed to robustly and comprehensively answer users’ queries on their captured memories. OmniQuery has two key components: (i) a question-agnostic pipeline to augment captured memories with contextual information extracted from other related memories to produce context-augmented memories, and (ii) a natural language question answering system that retrieves these processed memories and generates comprehensive answers with referenced captured memories as evidence. The processing pipeline in (i) is informed by the taxonomy of contextual information that we generated from a one-month diary study with 29 participants. Specifically, we collected and analyzed 299 user queries, identifying the types of contextual information that would be most useful for augmenting capture memories. We identified three types of personal questions (direct content queries, contextual filters, and hybrid queries) and generated a taxonomy of three major categories (atomic context, composite context, and semantic knowledge) which then guide the design of memory augmentation pipeline to enhance memory retrieval comprehensiveness. For (ii), OmniQuery employs a retrieval-augmented architecture: given a user input query, it (1) augments the query via a rewriting strategy, and (2) retrieves related memories from the augmented data and uses an LLM to generate the final answer, along with the reference memory instances.
@@
To evaluate OmniQuery, we conducted a user evaluation with 10 participants against a generic RAG-based baseline. The participants tested queries both logged during the diary study and generated during the evaluation session on a subset of their own captured memories. For each tested query, participants rated the user perceived correctness and completeness of the answers generated by both systems in a blinded manner. The results show that OmniQuery effectively answers different types of queries on users’ personal memories, outperforming the baseline with higher accuracy (71.5%, exceeding the baseline by 27.6%) and winning or tying 74.5% of the time in direct comparisons, which are based on user-perceived accuracy and completeness.
@@
  • A taxonomy of contextual information for augmenting captured memories, consisting of three major categories. The taxonomy was derived from authentic queries collected in a one-month diary study with 29 participants.
  • A pipeline of augmenting captured memories that leverages temporal-based reasoning to extract and infer missing contextual information from other related captured memories.
  • The design and implementation of an end-to-end system that comprehensively answers user queries.
  • A user evaluation of OmniQuery against a baseline system, showing OmniQuery’s effectiveness with 71.5% accuracy and outperforming the baseline (winning or tying 74.5% of the time).
@@ 2. Related Work
OmniQuery is inspired by and related to prior work in the areas on personal memory augmentation, multimodal question answering and applications that utilize contextual information.
@@ 2.1. Personal Memory Augmentation
A large body of work in human-computer interaction (HCI) has been focused on augmenting users’ memory. This includes developing reminder tools for elderlies or people with memory impairments (Caprani et al., 2006; Jamieson et al., 2014, 2017; Shin et al., 2021), providing proactive support in daily tasks (Zulfikar et al., 2024; Chan, 2022), or manipulating users’ memory focus in extended reality (Bonnail et al., 2023). These work typically focus on the “capturing” stage of the memory augmentation, where researchers develop wearable devices that continuously capture data using designated sensors, which record various modalities such as videos (Hodges et al., 2006; Dubourg et al., 2016; Mann, 1996), audios (Hayes et al., 2004; Vemuri et al., 2004), or bio-signals (Chan, 2020), to augment the memory database. For example, recent work such as Memoro developed a wearable, audio-based device that continuously records users’ conversations and enables memory suggestions in real-time, either through explicit queries or query-less contextual cues (Zulfikar et al., 2024). Differently, OmniQuery focuses on the “post-capturing” stage, utilizing already-existing memory data (e.g., photos and videos users have already captured). It addresses challenges in processing, annotating, and augmenting captured memories with contextual information.
Prior work in natural language processing (NLP), computer vision (CV), and information retrieval (IR) has studies methods of augmenting people memory. Perhaps the most related is QA on egocentric videos, which are also a form of personal data. Representative tasks include episodic memory retrieval (Gurrin et al., 2014; Grauman et al., 2021; Engel et al., 2023), where the system, given a long egocentric video and a query, localizes the answer within the video. However, these datasets differ from the captured data targeted by OmniQuery. The main challenge in egocentric videos is filtering through large, often noisy data, using data-driven approaches to train models for feature extraction. In contrast, captured memories represent a smaller, intentionally collected dataset, where the challenge lies in integrating scattered contextual information across multiple implicitly related memories. Therefore, OmniQuery employs a taxonomy-based method to augment existing data without the need for specific model training, improving QA performance.
@@ 2.2. Multimodal Question Answering
Over time, natural language QA research has shifted to more complex settings, including QA across different modalities (e.g., images (Agrawal et al., 2015; Goyal et al., 2016), videos (Yang et al., 2003; Zhang et al., 2023b; Maaz et al., 2023), tables (Zhang et al., 2023a) or knowledge graph (Yasunaga et al., 2021b; Huang et al., 2019b)), QA on large datasets (Lewis et al., 2021; Chen et al., 2017) and tasks that require multihop reasoning (Yang et al., 2018; Mavi et al., 2024). Recent advancements in large language models (LLMs) and multimodal foundation models (e.g., (Liu et al., 2024b, 2023a, 2023b)) have have enabled improved reasoning and answer generation over large, multimodal datasets. This is similar to OmniQuery’s use case as answering personal questions requires handling large amounts of captured memories and performing complex reasoning. Prior work has used retrieval-augmented generation (RAG) workflow (Lewis et al., 2021), which retrieve relevant information from external datasets based on a query and then generate output using the retrieved results. For example, MuRAG leverages RAG to answer open questions via retrieving related information from databases of images and text (Chen et al., 2022). VideoAgent leverages structured memories processed from long videos to accomplish video understanding tasks (Fan et al., 2024). However, these methods rely on datasets already rich in context (e.g., Wikipedia [^1] https://www.wikipedia.org/) and improvements are often achieved by designing new retrieval workflows (e.g., Self-RAG (Asai et al., 2023) and tree-based retrieval (Sarthi et al., 2024)) or query augmentation (Chan et al., 2024). More recently, GraphRAG introduced a data augmentation approach that generates a knowledge graph from extensive raw data to tackle tasks requiring higher-level understanding, such as query-focused summarization (Edge et al., 2024). While GraphRAG leverages data-driven methods to identify themes and communities within graph nodes, OmniQuery extends this concept by adopting a taxonomy-based augmentation approach, informed by insights from a diary study, to enhance retrieval on personal captured memories.
@@ 2.3. Applications Utilizing Contextual Information
Contextual information has long been important in HCI research from early mixed-initiative systems (Horvitz, 1999) to recent agentic workflow (Jaber et al., 2024). Over the past few years, there has been a surge in the usage of AI and LLMs in the HCI community, which enables extracting contextual information from processing raw multimodal information. For example, Li et al. studied how visually impaired people cook and emphasized the importance of conveying contextual information to users through multimodal models (Li et al., 2024a). Additionally, Human I/O leverages egocentric perceptions of users and detect situational impairments through reasoning on the multimodal sensing data (Liu et al., 2024a). GazePointAR develops a context-aware voice assistant to disambiguate users’ intent when interacting with real-world information (Lee et al., 2024). OmniActions categorizes digital follow-up actions on real-world information and provides proactive action prediction based on perceived context (Li et al., 2024b). Specifically, these system utilized off-the-shelf multimodal models to process raw sensory data, and leverages the reasoning capabilities of LLMs to infer the semantic context. OmniQuery builds on this approach by applying these AI techniques to extract and integrate semantic context scattered across various unstructured, raw captured memories. This augmentation enhances users’ memory databases, enabling them to answer personal questions about their memories through natural language queries.
3. DIARY STUDY
@@ 3.1. Method
To answer the research question above, we conducted a diary study, a methodology that enables participants to log data whenever need arose (Sohn et al., 2008). Specifically, we adopted the snippet-based technique proposed by Brandt et al. (Brandt et al., 2007). We asked participants to log queries on their past memories only when they had real intent under a genuine context, rather than brainstorming potential questions they might ask to retrieve specific past memories. This approach enabled us to collect authentic and spontaneous queries that users have in real-world scenarios.
We collect the data including: (i) the queries participants would use to retrieve or ask about their past memories, (ii) the reasons and contexts prompting these queries (e.g., wanting to show a past experience while chatting with a friend) and (iii) (optional) whether they were able to retrieve the corresponding memories from their album, and if so, how they did it (e.g., by scrolling through the photo album).
@@ 3.2. Participants
Thirty-two participants (i.e., 14 male, 17 female, one binary) were initially recruited through an online RSVP form distributed via the X platform [^2] https://x.com/. Participants come from North America and Asia. Eleven participants reported using Android devices, while the remainder used iOS devices in their daily life. Additionally, 16 participants reported actively logging their daily lives, 13 regularly logged important events and memorable experiences, 2 logged only essential information, and one seldom logged their lives. While participants were compensated based on their participation (\$50 for full participation), they were not required to log a specific number queries each day or over the entire study period. This approach was intentional, as we did not want to pressure them into generating queries artificially.
3.3. Data Summary and analysis
During the diary study, one participant opted out during the first week, and two participants did not log any queries throughout the entire study. Of the remaining participants, seven stopped logging after the first week. The rest remained active until the end of the study. As a result, we collected a total of 299 queries. On average, each participant contributed 10.27 queries, with a standard deviation of 6.09, and the highest number of queries contributed by a single participant was 25.
From the collected queries, we identified three major types of query: (1) direct content queries (75 queries), (2) context-based filters (28 queries), and (3) hybrid queries (191 queries). The rest five queries fell outside these categories as participants were attempting non memory-related tasks, such as “Mark yesterday pictures as favorites”.

### v2

@@ Abstract.
People often capture memories through photos, screenshots, and videos. While existing AI-based tools enable querying this data using natural language, they only support retrieving individual pieces of information like certain objects in photos, and struggle with answering more complex queries that involve interpreting interconnected memories like sequential events. We conducted a one-month diary study to collect realistic user queries and generated a taxonomy of necessary contextual information for integrating with captured memories. We then introduce OmniQuery, a novel system that is able to answer complex personal memory-related questions that require extracting and inferring contextual information. OmniQuery augments individual captured memories through integrating scattered contextual information from multiple interconnected memories. Given a question, OmniQuery retrieves relevant augmented memories and uses a large language model (LLM) to generate answers with references. In human evaluations, we show the effectiveness of OmniQuery with an accuracy of 71.5%, outperforming a conventional RAG system by winning or tying for 74.5% of the time.
@@
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713448isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Natural language interfacesccs: Human-centered computing User studies
Figure 1. OmniQuery is able to answer complex personal questions (a) on individuals’ captured memories (b), such as captured photos, saved screenshots, and recorded videos. It augments the captured memories by identifying and integrating contextual information scattered across multiple interconnected memories (c). OmniQuery then uses this information to retrieve relevant memories and leverages an LLM to generate a comprehensive answer with reference memories (d).
@@ 1. INTRODUCTION
People often record their everyday life by taking photos, screenshots, and videos for saving important information, documenting special occasions, or simply capturing a funny moment (Li et al., 2024b). These recorded instances, referred to as captured memories, collectively represent subsets of an individual’s episodic memories (Tulving, 2002), a type of long-term memory that contains both specific past experiences and associated contextual details. These episodic memories are essential for answering higher-level memory-related personal questions like ”What social events did I attend during CHI 2024?” (Figure 1a). Being able to do so could help users reflect on past experiences and make informed decisions in daily tasks.
@@
However, these raw captured memories by themselves are insufficient to answer personal questions, as they lack contextual details that are typically implicit and scattered across multiple pieces of data. As shown in Figure 1b, memories of attending parties during CHI 2024 are not explicitly annotated as occurring during the event. Answering such personal questions requires extracting and integrating contextual information not contained within a single captured instance. For example, by integrating multiple memories that mention “CHI 2024” in their content and extracting their metadata, it is possible to determine when the users attended the conference and connect related social events memories from that period to CHI 2024 (Figure 1c), enabling the answering of the query (Figure 1d).
@@
Advancements in AI have enable question answering (QA) on long documents (Beltagy et al., 2020; Wang et al., 2024), knowledge graph (Huang et al., 2019a; Yasunaga et al., 2021a), multimodal databases (Talmor et al., 2021; Chen et al., 2022), and egocentric videos (Mangalam et al., 2023; Hou et al., 2023). These methods typically rely on data-driven approaches to train powerful models for the target task. However, the private nature of captured memories makes it difficult to curate large datasets, posing challenges for training models specifically for QA on personal data. Recent LLM-based work has adopted retrieval augmented generation (RAG) workflows to handle external databases without specific training (Lewis et al., 2021). However, such methods depend on explicit connections between queries and relevant external data (Edge et al., 2024). In contrast, captured memories are often unstructured and lack contextual annotations, making it difficult to establish explicit links between queries and scattered memories.
@@
To facilitate QA on personal captured memories, we propose OmniQuery, a novel approach designed to robustly and comprehensively answer users’ queries on their captured memories. OmniQuery has two key components: (i) a question-agnostic pipeline to augment captured memories with contextual information extracted from other related memories to produce context-augmented memories, and (ii) a natural language QA system that retrieves these processed memories and generates comprehensive answers with referenced captured memories as evidence. The design of OmniQuery is informed by a taxonomy of contextual information that we generated from a one-month diary study with 29 participants. Specifically, we collected and analyzed 299 user queries to identify three types of personal questions (direct content queries, contextual filters, and hybrid queries) and three categories of contextual information (atomic context, composite context, and semantic knowledge). For (ii) , OmniQuery employs a retrieval-augmented architecture: given a user input query, it augments the query via a rewriting strategy, retrieves related memories from the augmented data, and generates the final answer with referenced memory instances via an LLM.
@@
To evaluate OmniQuery, we conducted a user evaluation with 10 participants against a generic RAG-based baseline. The participants tested queries both logged during the diary study and generated during the evaluation session on a subset of their own captured memories. For each tested query, participants rated the user perceived correctness and completeness of the answers generated by both systems in a blinded manner. The results show that OmniQuery effectively answers different types of queries on users’ personal memories, outperforming the baseline with higher accuracy (71.5%, exceeding the baseline by 27.6%) and winning or tying 74.5% of the time in direct comparisons.
@@
  • A taxonomy of contextual information for augmenting captured memories, derived from queries collected in a one-month diary study with 29 participants.
  • A taxonomy-based pipeline of augmenting captured memories that leverages temporal-based reasoning to extract and infer missing contextual information from other related memories.
  • The design and implementation of an end-to-end taxonomy-informed system for personal QA [^1] OmniQuery is open-source at: https://github.com/ljhnick/omniquery .
  • A user evaluation of OmniQuery against a baseline system, demonstrating OmniQuery’s effectiveness with 71.5% accuracy and outperforming the baseline (winning or tying 74.5% of the time).
@@ 2.1. Personal Memory Augmentation
A large body of work in human-computer interaction (HCI) has explored how to augment users’ memories. This includes developing reminder tools for elderlies or people with memory impairments (Caprani et al., 2006; Jamieson et al., 2014, 2017; Shin et al., 2021), providing proactive support in daily tasks (Zulfikar et al., 2024; Chan, 2022), or manipulating users’ memory focus in extended reality (Bonnail et al., 2023). These works typically focus on the “capturing” stage of the memory augmentation, where researchers develop wearable devices that continuously capture data using designated sensors, which record various modalities such as videos (Hodges et al., 2006; Dubourg et al., 2016; Mann, 1996), audios (Hayes et al., 2004; Vemuri et al., 2004), or bio-signals (Chan, 2020), to augment the memory database. For example, recent work such as Memoro developed a wearable, audio-based device that continuously records users’ conversations and enables memory suggestions in real-time, either through explicit queries or query-less contextual cues (Zulfikar et al., 2024). Differently, OmniQuery focuses on the “post-capturing” stage, utilizing already-existing memory data (e.g., photos and videos users have already captured). It addresses challenges in processing, annotating, and augmenting captured memories with contextual information.
Prior work in natural language processing (NLP), computer vision (CV), and information retrieval (IR) has studied methods of augmenting people’s memory. Perhaps the most related is QA on egocentric videos, which are also a form of personal data. Representative tasks include episodic memory retrieval (Gurrin et al., 2014; Grauman et al., 2021; Engel et al., 2023), where the system, given a long egocentric video and a query, localizes the answer within the video. However, these datasets differ from the captured data targeted by OmniQuery. The main challenge in egocentric videos is filtering through large, often noisy data, using data-driven approaches to train models for feature extraction. In contrast, captured memories represent a smaller, intentionally collected dataset, where the challenge lies in integrating scattered contextual information across multiple implicitly related memories. Therefore, OmniQuery employs a taxonomy-based method to augment existing data without the need for specific model training, improving QA performance.
@@ 2.2. Multimodal Question Answering
Over time, natural language QA research has shifted to more complex settings, including QA across different modalities (e.g., images (Agrawal et al., 2015; Goyal et al., 2016), videos (Yang et al., 2003; Zhang et al., 2023b; Maaz et al., 2023), tables (Zhang et al., 2023a) or knowledge graph (Yasunaga et al., 2021b; Huang et al., 2019b)), QA on large datasets (Lewis et al., 2021; Chen et al., 2017) and tasks that require multi-hop reasoning (Yang et al., 2018; Mavi et al., 2024). Recent advancements in large language models (LLMs) and multimodal foundation models (e.g., (Liu et al., 2024b, 2023a, 2023b)) have enabled improved reasoning and answer generation over large, multimodal datasets. This is similar to OmniQuery’s use case as answering personal questions requires handling large amounts of captured memories and performing complex reasoning. Prior work has used retrieval-augmented generation (RAG) workflow (Lewis et al., 2021), which retrieves relevant information from external datasets based on a query and then generates output using the retrieved results. For example, MuRAG leverages RAG to answer open questions via retrieving related information from databases of images and text (Chen et al., 2022). VideoAgent leverages structured memories processed from long videos to accomplish video understanding tasks (Fan et al., 2024). However, these methods rely on datasets already rich in context (e.g., Wikipedia [^2] https://www.wikipedia.org/) and improvements are often achieved by designing new query augmentation (Chan et al., 2024) and retrieval workflows such as Self-RAG (Asai et al., 2023) and tree-based retrieval (Sarthi et al., 2024).
More recently, GraphRAG introduced a data augmentation approach that extracts a knowledge graph from raw data to tackle tasks requiring higher-level understanding, such as query-focused summarization (Edge et al., 2024). While we do not explicitly employ a graph data structure in OmniQuery, we adopt GraphRAG’s structured, hierarchical approach for RAG-based tasks and extend it with taxonomy-based augmentation informed by insights from a diary study to enhance retrieval results on personal captured memories. Finally, when it comes to QA system design, Jim Gray proposed the “20 queries” heuristic that optimizes for answering a core set of questions to address the long tail distribution of potential queries (Szalay, 2008). We adopt the same design principle and replace the specific rules with our contextual information taxonomy.
@@ 2.3. Applications Utilizing Contextual Information
Contextual information has long been important in HCI research from early mixed-initiative systems (Horvitz, 1999) to recent agentic workflows (Jaber et al., 2024). Over the past few years, there has been a surge in the usage of AI and LLMs in the HCI community to extract contextual information from processing raw multimodal information. For example, Li et al. studied how visually impaired people cook and emphasized the importance of conveying contextual information to users through multimodal models (Li et al., 2024a). Additionally, Human I/O leverages egocentric perceptions of users and detect situational impairments through reasoning on the multimodal sensing data (Liu et al., 2024a). GazePointAR develops a context-aware voice assistant to disambiguate users’ intent when interacting with real-world information (Lee et al., 2024). OmniActions categorizes digital follow-up actions on real-world information and provides proactive action prediction based on perceived context (Li et al., 2024b). These system utilized off-the-shelf multimodal models to process raw sensory data and leverage the reasoning capabilities of LLMs to infer the semantic context. OmniQuery builds on this approach by applying these AI techniques to extract and integrate semantic context scattered across various unstructured, raw captured memories. This augmentation enhances users’ memory databases, enabling them to answer personal questions about their memories through natural language queries.
3. DIARY STUDY: UNDERSTANDING USER QUERIES
@@ 3.1. Method
To answer the research question above, we conducted a diary study, a methodology that enables participants to log data whenever need arose (Sohn et al., 2008). Specifically, we adopted the snippet-based technique proposed by Brandt et al. (Brandt et al., 2007). We asked participants to log queries on their past memories only when they had real intent under a genuine context, rather than brainstorming potential questions they might ask to retrieve specific past memories. This approach enabled us to collect spontaneous, authentic queries that users have in real-world scenarios.
We collect the data including: (i) the queries participants would use to retrieve or ask about their past memories, (ii) the reasons and contexts of these queries (e.g., wanting to show a past experience while chatting with a friend) and (iii) (optional) whether they were able to retrieve the corresponding memories from their album, and if so, how they did it (e.g., by scrolling through the photo album).
@@ 3.2. Participants
32 participants (14 male, 17 female, and one non-binary) were initially recruited through an online RSVP form distributed via the X platform [^3] https://x.com/. Participants came from North America and Asia. 11 participants reported using Android devices, while the remainder used iOS devices in their daily life. Additionally, 16 participants reported actively logging their daily lives, 13 regularly logged important events and memorable experiences, two logged only essential information, and one seldom logged their lives. While participants were compensated based on their participation (\$50 for full participation), they were not required to log a specific number queries each day or over the entire study period. This approach was intentional, as we did not want to require them to generating queries artificially.
3.3. Data Summary
During the diary study, one participant opted out during the first week, and two participants did not log any queries. Of the remaining participants, seven stopped logging after the first week. The rest remained active until the end of the study. As a result, we collected a total of 299 queries. On average, each participant contributed 10.27 queries ($\mathrm{SD}=6.09$). The highest number of queries from a single participant was 25 and the lowest was 3.
From the collected queries, we identified three types of query: (1) direct content queries (75 queries), (2) context-based filters (28 queries), and (3) hybrid queries (191 queries). The remaining five queries fell outside of these categories as participants attempted non-memory-related tasks like“Mark yesterday pictures as favorites”.